# Volumetric Functional Maps

### v2

@@ 4. Method
The functional maps framework is not bound to the domain of 2-dimensional surfaces. Indeed, the Laplace-Beltrami Operator (LBO) can be defined on Riemannian manifolds of arbitrary dimension and it always admits a spectral decomposition. We discretize the LBO of a tetrahedral mesh $\mathcal{M}=(V_{\mathcal{M}},T_{\mathcal{M}})$ using the $n$-dimensional cotangent formula (Liao et al., 2009; Crane, 2019) which defines the discrete LBO as a pair of matrices $\mathbf{S}$ and $\mathbf{W}$, respectively denoted as stiffness and mass matrices. Its spectral decomposition can be obtained by solving the generalized eigenproblem $\mathbf{S}\Phi=\mathbf{W}\Phi\Lambda$.
@@
We provide in Appendix A the implementation details, as well as a brief discussion about the choice of the LBO and its discretization. However, we highlight that, since most applications that we address involve analyzing information at the surface (which is the boundary of a tetrahedral mesh), we are forced to use a discretization that imposes Neumann’s boundary conditions to avoid zero-valued eigenfunctions on the surface.
4.1. Functional Maps for Tetrahedral Meshes
@@
Most successful algorithms for the estimation of the functional map between shapes rely on the property of intrinsic structures and processes to behave similarly on similar surfaces. For instance, the invariance of the LBO basis and the heat diffusion under isometric and quasi-isometric deformations plays a key role in stable functional map estimation methods (Nogneng and Ovsjanikov, 2017; Nogneng et al., 2018; Ovsjanikov et al., 2016). We provide a discussion on isometric deformations between volumes in Appendix C.
@@
Figure 2. Top: the first 5 non-constant LBO eigenfunctions on two humanoid non-isometric shapes. Bottom: eigenvalue errors distribution for the volumetric LBO (blue) and the surface LBO (orange).
LBO Eigenbasis.
When dealing with continuous Riemannian manifolds, the LBO and its eigenbasis are invariant to isometric deformations. However, in the discrete setting the position and connectivity of the vertices have a significant impact on the equivalence of the operator across different meshes, and changing the triangulation of a surface or its vertex density can strongly affect the spectral decomposition (Lescoat et al., 2020).
Moving from discrete surfaces to discrete volumes makes this problem even more evident. In a regular triangular mesh, each vertex has on average 6 neighboring vertices and 6 neighboring triangles, while in regular tetrahedral meshes the average size of a vertex neighborhood is 12 vertices and 20 tetrahedra.
To ensure that the LBO has a consistent behavior across isometric or similar volumetric shapes, we compute the spectral decomposition of the LBO on 40 pairs of shapes with identical connectivity released by Su et al. (2019) and compare both the eigenvalues and the eigenfunctions of the LBO. Our results are summarized in Figure 2. The first two rows of the figure show that two humanoid shapes share the same LBO eigenfunctions (up to sign flips), even if they are not isometric and not representing the same subject: the overall similarity is sufficient to generate the same low-frequency spectrum. To compare the eigenvalues, we first compute the vectors $\bm{\lambda}_{\mathcal{M}},\bm{\lambda}_{\mathcal{N}}$ of the first 100 LBO eigenvalues on two tetrahedral meshes. Since the absolute difference between LBO eigenvalues is notoriously not informative, we rely on both the relative difference and the offset difference proposed by Moschella et al. (2022). For a deeper discussion about the error measures, we refer to Appendix B.
For each pair of tetrahedral meshes, we also extract the triangular meshes representing their boundary surface and compute the same quantities. For a more meaningful comparison, all the tetrahedral meshes are rescaled to have unitary volume, and all the surface meshes are rescaled to have unitary surface area. The error distributions are shown at the bottom of Figure 2, where we can see that the spectral similarity between pairs of tetrahedral and surface meshes is comparable.
Figure 3. Distribution of the orthogonality (left) and Laplacian commutativity (right) metrics for pairs of volumetric shapes (blue) and surface meshes (orange).
Volumetric Functional Maps.
Given a definition for the discrete LBO, the definition of the functional maps framework in tetrahedral meshes follows directly from its 2-dimensional counterpart (see Section 3). Nonetheless, in order to ensure a seamless integration of the existing approaches in the volumetric setting, we study how the usual properties of a functional map behave on tetrahedral meshes.
Given two meshes $\mathcal{M}$ and $\mathcal{N}$ and a functional map $\mathbf{C}:\mathcal{F}(\mathcal{M})\to\mathcal{F}(\mathcal{N})$, Lescoat et al. (2020) propose two metrics for evaluating the ideal qualities of $\mathbf{C}$: the orthogonality $\|\mathbf{C}\|_{O}$ and the Laplacian commutativity $\|\mathbf{C}\|_{L}$
(5) $\|\mathbf{C}\|_{O}=\frac{\|\mathbf{C}^{\top}\mathbf{C}-\mathbf{I}\|}{\|\mathbf{I}\|}\,,\qquad\|\mathbf{C}\|_{L}=\frac{\|\mathbf{C}\Lambda_{\mathcal{M}}-\Lambda_{\mathcal{N}}\mathbf{C}\|}{\|\mathbf{C}\Lambda_{\mathcal{M}}\|}\,,$
@@
Again, we compute the same quantities for tetrahedral and triangular meshes across the same 40 pairs of shapes, summarizing the results in Figure 3. As shown in the figure, the quality of volumetric functional maps is comparable to the surface cases.
4.2. Spectral Connectivity Transfer
The success of the functional maps framework is also due to its ability of smoothing out signals and correspondences. Indeed, the use of a truncated basis results in filtering out high-frequency details and producing a smoothly interpolated signal.
This smoothness property can be combined with the fact that the basis is usually much smaller than the number of vertices. As a result, the spectral representation of a signal can be approximated even if the signal is noisy or partially absent from some regions of the shape. This idea has been exploited by Melzi et al. (2019b) for efficiently approximating a functional map using only a reduced sampling of the original surface. In our setting, we take advantage of this property for transferring the connectivity between volumes.
Functional Maps for Connectivity Transfer.
Let $\mathcal{M}=(V_{\mathcal{M}},T_{\mathcal{M}})$ be a tetrahedral mesh and $\partial\mathcal{M}=(V_{\partial\mathcal{M}},F_{\partial\mathcal{M}})$ be its external surface. Let us consider another surface $\partial\mathcal{N}=(V_{\partial\mathcal{N}},F_{\partial\mathcal{N}})$ in known correspondence with $\partial\mathcal{M}$ through the (surface) mapping function $\pi:\partial\mathcal{M}\to\partial\mathcal{N}$. There are various options in the literature to obtain the tetrahedralization $\mathcal{N}$ of the interior of $\partial\mathcal{N}$ (Hang, 2015; Diazzi et al., 2023). However, by no mean the connectivity generated by these algorithms will match the connectivity of the mesh $\mathcal{M}$.
Given the surface $\partial\mathcal{N}$ and its induced tetrahedral mesh $\mathcal{N}$, we exploit the known correspondence $\pi:\partial\mathcal{M}\to\partial\mathcal{N}$ to induce a full correspondence $\pi^{\prime}:\mathcal{M}\to\mathcal{N}$. Let $\Phi_{\mathcal{M}},\Phi_{\mathcal{N}}$ be the LBO eigenbases of the volumes $\mathcal{M}$ and $\mathcal{N}$, respectively, and let us denote with $\Phi_{\partial\mathcal{M}}=(\Phi_{\mathcal{M}})_{\partial},\Phi_{\partial\mathcal{N}}=(\Phi_{\mathcal{N}})_{\partial}$ their restrictions to the surfaces $\partial\mathcal{M}$ and $\partial\mathcal{N}$, respectively. We stress that $\Phi_{\partial\mathcal{M}}$ is not the LBO eigenbasis of $\partial\mathcal{M}$. We note that the functional map $\mathbf{C}$ induced by $\pi^{\prime}$ must be such that
(6) $\Phi_{\mathcal{M}}\mathbf{C}=T_{\pi^{\prime}}(\Phi_{\mathcal{N}})\,.$
(7) $\Phi_{\partial\mathcal{M}}\mathbf{C}=(\Phi_{\mathcal{M}})_{\partial}\mathbf{C}=(\Phi_{\mathcal{M}}\mathbf{C})_{\partial}=(T_{\pi^{\prime}}(\Phi_{\mathcal{N}}))_{\partial}=T_{\pi}(\Phi_{\partial\mathcal{N}})\,.$
In general, the size of the basis (i.e., the number of columns of $\Phi_{\partial\mathcal{M}}$) is significantly smaller than the number of vertices (i.e., the number of rows), even if we only consider the vertices at the surface. Intuitively, this suggests that the restricted eigenfunctions $\Phi_{\partial\mathcal{M}}$ at the surface are linearly independent. Indeed, classical results from control theory support this intuition, proving that the boundary traces (i.e., the restriction to the boundary) of the eigenfunctions of any elliptic operator are linearly independent in a system with Neumann boundary conditions (Lasiecka and Triggiani, 1983; Triggiani, 2008) (see Appendix D for a deeper analysis of the descriptive power of the boundary traces). Therefore, $\Phi_{\partial\mathcal{M}}$ admits left inverse and the volumetric functional map can be approximated via the known surface correspondence as $\mathbf{C}\approx\Phi_{\partial\mathcal{M}}^{\dagger}T_{\pi}(\Phi_{\partial\mathcal{N}})$.
(8) $\begin{gathered}T_{\pi^{\prime}}(x_{\mathcal{N}})=\Phi_{\mathcal{M}}\mathbf{C}\Phi_{\mathcal{N}}^{\dagger}x_{\mathcal{N}}\,,\\ T_{\pi^{\prime}}(y_{\mathcal{N}})=\Phi_{\mathcal{M}}\mathbf{C}\Phi_{\mathcal{N}}^{\dagger}y_{\mathcal{N}}\,,\\ T_{\pi^{\prime}}(z_{\mathcal{N}})=\Phi_{\mathcal{M}}\mathbf{C}\Phi_{\mathcal{N}}^{\dagger}z_{\mathcal{N}}\,.\end{gathered}$
@@ 4.2. Spectral Connectivity Transfer
By using the values $T_{\pi^{\prime}}(x_{\mathcal{N}}),T_{\pi^{\prime}}(y_{\mathcal{N}}),T_{\pi^{\prime}}(z_{\mathcal{N}})$ as coordinates for the vertices of $\mathcal{M}$, we get a connectivity transfer from $\mathcal{M}$ onto $\mathcal{N}$.
@@
Figure 4. A visual representation of the pipeline for extrapolating the interior coordinates from the surface correspondence. Given the surface correspondence $\pi$ (first row), we approximate the spectral embedding of the surface coordinates of $\partial\mathcal{N}$ using the boundary restriction of the eigenfunctions of $\mathcal{M}$ (second row). Then, we reconstruct the interior coordinates from the spectral embedding using the eigenfunctions on the entire volume (third row), and we use the reconstructed coordinates to transfer the interior connectivity (fourth row).
Spectral Representation of Volumes.
@@
Intuitively, a watertight surface already encodes information about the enclosed volume. Since the LBO eigenbasis provides smoothness guarantees in signal reconstruction, it is a good candidate for extrapolating the interior coordinates of a volume from its surface.
@@
Knowing the ground truth correspondence $\pi:\partial\mathcal{M}\to\partial\mathcal{N}$, we can transfer the surface coordinates $x_{\partial\mathcal{N}},y_{\partial\mathcal{N}},z_{\partial\mathcal{N}}$ of $\partial\mathcal{N}$ onto $\partial\mathcal{M}$ as $T_{\pi}(x_{\partial\mathcal{N}}),T_{\pi}(y_{\partial\mathcal{N}}),T_{\pi}(z_{\partial\mathcal{N}})$. We then consider the LBO eigenbasis $\Phi_{\mathcal{M}}$ and its restriction $\Phi_{\partial\mathcal{M}}$ to the surface. Following a reasoning analogous to the discussion in the previous paragraph, we have that
(9) $\begin{gathered}(T_{\pi^{\prime}}(x_{\mathcal{N}}))_{\partial}\approx(\Phi_{\mathcal{M}}\Phi_{\mathcal{M}}^{\dagger}T_{\pi^{\prime}}(x_{\mathcal{N}}))_{\partial}=\Phi_{\partial\mathcal{M}}\Phi_{\mathcal{M}}^{\dagger}T_{\pi^{\prime}}(x_{\mathcal{N}})\,,\\ T_{\pi}(x_{\partial\mathcal{N}})\approx\Phi_{\partial\mathcal{M}}\Phi_{\partial\mathcal{M}}^{\dagger}T_{\pi}(x_{\partial\mathcal{N}})\,.\end{gathered}$
@@
(10) $\begin{gathered}T_{\pi^{\prime}}(x_{\mathcal{N}})=\Phi_{\mathcal{M}}\Phi_{\partial\mathcal{M}}^{\dagger}T_{\pi}(x_{\partial\mathcal{N}})\,,\\ T_{\pi^{\prime}}(y_{\mathcal{N}})=\Phi_{\mathcal{M}}\Phi_{\partial\mathcal{M}}^{\dagger}T_{\pi}(y_{\partial\mathcal{N}})\,,\\ T_{\pi^{\prime}}(z_{\mathcal{N}})=\Phi_{\mathcal{M}}\Phi_{\partial\mathcal{M}}^{\dagger}T_{\pi}(z_{\partial\mathcal{N}})\,.\end{gathered}$
A visual representation of our pipeline is shown in Figure 4.
@@
Concrete examples of reliable connectivity transfer using both a functional transfer of interior coordinates and their extrapolation from the spectral embedding of the surface coordinates are depicted in Figure 9 (resp. Transfer and Extrapolation). For this specific test, we extended the LBO eigenbasis with the Coordinates Manifold Harmonics (CMH) basis (Marin et al., 2019), which also includes three additional functions encoding the extrinsic per vertex $xyz$ coordinates and is known to be better suited to reconstruct extrinsic information. Additional experiments on the whole dataset, considering both the LBO and CMH eigenbasis are reported in Section 5.2.
4.3. Volume-Aware Surface Correspondence
@@
Besides purely volumetric applications, we observe that a good functional map among volumes also induces a good functional map among their external surfaces. Intuitively, a volumetric shape is a richer structure than a surface, yielding more information both locally and globally. Therefore, we guess that functional maps among tetrahedral meshes are more accurate and informative than maps among triangular meshes. This idea is grounded from the results of Raviv et al. (2010), who have shown that using volume-informed shape descriptors has a positive effects in computing correspondences between surfaces.
As mentioned in Section 4.2, the functional map that relates the LBO eigenbases of two volumetric shapes also relates their restriction to the surfaces. As a consequence, the volumetric map can be used to infer a correspondence between surfaces.
Given two surface meshes $\partial\mathcal{M},\partial\mathcal{N}$, we compute the tetrahedral meshes $\mathcal{M},\mathcal{N}$ representing their interior volume. In general, the tetrahedralization of a surface does not always produce a volume with the same surface connectivity as the original mesh, but if the two shapes can be closely aligned in 3D the scalar functions can be faithfully transferred using nearest surface point projections with barycentric interpolation (Maggioli et al., 2025).
@@
Once we have the tetrahedral meshes $\mathcal{M},\mathcal{N}$, we can compute the functional map $\mathbf{C}$ relating their eigenbases via $\Phi_{\mathcal{M}}\mathbf{C}=T_{\pi}(\Phi_{\mathcal{N}})$, and we can use that same functional map for defining the surface relation $(\Phi_{\mathcal{M}})_{\partial}\mathbf{C}=(T_{\pi}(\Phi_{\mathcal{N}}))_{\partial}$. Therefore, the surface correspondence can be extracted via the nearest neighbor search $\mathrm{NN}(\Phi_{\partial\mathcal{M}}\mathbf{C},\Phi_{\partial\mathcal{N}})$.
5. Results and Applications
We implemented our software pipeline in Matlab, starting from the reference implementations for the functional map estimation with descriptors preservation (Nogneng and Ovsjanikov, 2017) and ZoomOut (Melzi et al., 2019b), extending these tools to enable volume spectral processing, as detailed in Section 4 and in the remainder of this section. Our code is available at anonymous.
5.1. Volumetric Functional Maps

### v3

@@ 4 Method
The functional maps framework is not bound to the domain of 2-dimensional surfaces. Indeed, the Laplace-Beltrami Operator (LBO) can be defined on Riemannian manifolds of arbitrary dimension, and it always admits a spectral decomposition. We discretize the LBO of a tetrahedral mesh $\mathcal{M}=(V_{\mathcal{M}},T_{\mathcal{M}})$ using the $n$-dimensional cotangent formula [57, 21] which defines the discrete LBO as a pair of matrices $\mathbf{S}$ and $\mathbf{W}$, denoted as stiffness and mass matrices, respectively. Its spectral decomposition can be obtained by solving the generalized eigenproblem $\mathbf{S}\Phi=\mathbf{W}\Phi\Lambda$.
@@
Implementation details and a discussion on the choice of the LBO are provided in Appendix A. We highlight that most applications we address involve analyzing information at the surface (i.e., the boundary of a tetrahedral mesh), compelling us to impose Neumann’s boundary conditions to prevent zero-valued eigenfunctions on the surface.
@@
Most successful algorithms for estimating the functional map between shapes rely on the property that intrinsic structures and processes behave similarly on similar surfaces. For instance, the invariance of the LBO basis and the heat diffusion under (quasi-)isometric deformations plays a key role in stable functional map estimation [86, 85, 88]. Despite this invariance holds for continuous volumetric shapes (see Fig. 2), in a discrete setting the position and connectivity of the vertices can heavily impact the numerical stability of the Laplacian operator [53]. Appendix B provides a detailed empirical study on the invariance of the LBO in the volumetric setting, and Appendix C contains a deeper discussion on isometric deformations between volumes.
@@
Figure 2: The first 5 non-constant volumetric LBO eigenfunctions on two humanoid non-isometric shapes
4.1 Spectral Connectivity Transfer
The success of functional mapping is also due to its ability to smooth out signals and correspondences: a truncated basis filters out high-frequency details, producing smoothly interpolated signals. This smoothness, combined with the fact that the basis is usually much smaller than the number of vertices, allows to approximate the spectral representation of noisy and partial signals. This feature has been exploited by Melzi et al. [79] to efficiently approximate a functional map using only a reduced sampling of the original surface. In our setting, we take advantage of this property to transfer connectivity between volumes.
Functional Connectivity Transfer.
Let $\mathcal{M}=(V_{\mathcal{M}},T_{\mathcal{M}})$ be a tetrahedral mesh with surface $\partial\mathcal{M}=(V_{\partial\mathcal{M}},F_{\partial\mathcal{M}})$. Let $\partial\mathcal{N}=(V_{\partial\mathcal{N}},F_{\partial\mathcal{N}})$ be another surface in known correspondence with $\partial\mathcal{M}$ through the (surface) map $\pi:\partial\mathcal{M}\to\partial\mathcal{N}$. There are various options to obtain a tetrahedralization $\mathcal{N}$ of the interior of $\partial\mathcal{N}$ [39, 24]. However, by no means, its connectivity will match the connectivity of $\mathcal{M}$.
Given a surface $\partial\mathcal{N}$ with induced tetrahedral mesh $\mathcal{N}$, we exploit the known correspondence $\pi:\partial\mathcal{M}\to\partial\mathcal{N}$ to induce a volumetric correspondence $\pi^{\prime}:\mathcal{M}\to\mathcal{N}$. Let $\Phi_{\mathcal{M}},\Phi_{\mathcal{N}}$ be the LBO eigenbases of the volumes $\mathcal{M}$ and $\mathcal{N}$, respectively, and let us denote with $\Phi_{\partial\mathcal{M}}=(\Phi_{\mathcal{M}})_{\partial},\Phi_{\partial\mathcal{N}}=(\Phi_{\mathcal{N}})_{\partial}$ their restrictions to the surfaces. We stress that $\Phi_{\partial\mathcal{M}}$ is not the LBO eigenbasis of $\partial\mathcal{M}$. The functional map $\mathbf{C}$ induced by $\pi^{\prime}$ must be such that
$\Phi_{\mathcal{M}}\mathbf{C}=T_{\pi^{\prime}}(\Phi_{\mathcal{N}})\,.$ (5)
$\Phi_{\partial\mathcal{M}}\mathbf{C}=(\Phi_{\mathcal{M}}\mathbf{C})_{\partial}=(T_{\pi^{\prime}}(\Phi_{\mathcal{N}}))_{\partial}=T_{\pi}(\Phi_{\partial\mathcal{N}})\,.$ (6)
In general, the size of the basis (i.e., the number of columns of $\Phi_{\partial\mathcal{M}}$) is significantly smaller than the number of vertices (i.e., the number of rows), even if we only consider the vertices at the surface. Intuitively, this suggests that the restricted eigenfunctions $\Phi_{\partial\mathcal{M}}$ at the surface are linearly independent. Indeed, classical results from control theory support this intuition, proving that the boundary traces (i.e., the restriction to the boundary) of the eigenfunctions of any elliptic operator are linearly independent in a system with Neumann boundary conditions [51, 117] (see the Appendix D for a deeper analysis of the descriptive power of the boundary traces). Therefore, $\Phi_{\partial\mathcal{M}}$ admits left inverse and the volumetric functional map can be approximated via the known surface correspondence as $\mathbf{C}\approx\Phi_{\partial\mathcal{M}}^{\dagger}T_{\pi}(\Phi_{\partial\mathcal{N}})$.
@@ 4.1 Spectral Connectivity Transfer
and similarly for $y_{\mathcal{N}}$ and $z_{\mathcal{N}}$. By using the values $T_{\pi^{\prime}}(x_{\mathcal{N}}),T_{\pi^{\prime}}(y_{\mathcal{N}}),T_{\pi^{\prime}}(z_{\mathcal{N}})$ as coordinates for the vertices of $\mathcal{M}$, we get a connectivity transfer from $\mathcal{M}$ onto $\mathcal{N}$.
@@
Figure 3: Our pipeline for extrapolating interior coordinates from surface correspondences. Given the surface map $\pi$ (first row), we approximate the spectral embedding of the surface coordinates of $\partial\mathcal{N}$ using the boundary restriction of the eigenfunctions of $\mathcal{M}$ (second row). Using the eigenfunctions on the entire volume, we reconstruct the interior coordinates from the spectral embedding (third row), and use these coordinates to transfer the inner connectivity (fourth row).
Spectral Extrapolation of Coordinates.
@@
A watertight surface already encodes information about the enclosed volume. Since the LBO provides smoothness guarantees in signal reconstruction, it is a good candidate for extrapolating the interior coordinates of a volume from its surface.
@@
Given the ground truth correspondence $\pi:\partial\mathcal{M}\to\partial\mathcal{N}$, we can transfer the surface coordinates $x_{\partial\mathcal{N}},y_{\partial\mathcal{N}},z_{\partial\mathcal{N}}$ of $\partial\mathcal{N}$ onto $\partial\mathcal{M}$ as $T_{\pi}(x_{\partial\mathcal{N}}),T_{\pi}(y_{\partial\mathcal{N}}),T_{\pi}(z_{\partial\mathcal{N}})$. We then consider the LBO eigenbasis $\Phi_{\mathcal{M}}$ and its restriction $\Phi_{\partial\mathcal{M}}$ to the surface. Following a reasoning analogous to the discussion in the previous paragraph, we have that
$\begin{gathered}(T_{\pi^{\prime}}(x_{\mathcal{N}}))_{\partial}\approx(\Phi_{\mathcal{M}}\Phi_{\mathcal{M}}^{\dagger}T_{\pi^{\prime}}(x_{\mathcal{N}}))_{\partial}=\Phi_{\partial\mathcal{M}}\Phi_{\mathcal{M}}^{\dagger}T_{\pi^{\prime}}(x_{\mathcal{N}})\,,\\ T_{\pi}(x_{\partial\mathcal{N}})\approx\Phi_{\partial\mathcal{M}}\Phi_{\partial\mathcal{M}}^{\dagger}T_{\pi}(x_{\partial\mathcal{N}})\,.\end{gathered}$ (8)
@@
[…]
and analogously for $y_{\mathcal{N}}$ and $z_{\mathcal{N}}$. A visual representation of our pipeline is shown in Fig. 3.
@@
Concrete examples of reliable connectivity transfer using both a functional transfer of interior coordinates and their extrapolation from the spectral embedding of the surface coordinates are depicted in Fig. 7 (resp. Transfer and Extrapolation). For this specific test, we extended the LBO eigenbasis with the Coordinates Manifold Harmonics (CMH) basis [73], which also includes three additional functions encoding the extrinsic per vertex xyzxyz coordinates and is known to be better suited to reconstruct extrinsic information. Additional experiments on the whole dataset, considering both the LBO and CMH eigenbasis are reported in Sec. 5.2.
4.2 Volume-Aware Surface Correspondence
@@
Besides purely volumetric applications, we observe that a good functional map among volumes also induces a good functional map among their external surfaces. Intuitively, a volumetric shape is a richer structure than a surface, yielding more information both locally and globally. Therefore, we guess that functional maps among tetrahedral meshes are more accurate and informative than maps among triangular meshes. This idea is grounded from the results of Raviv et al. [97], who have shown that using volume-informed shape descriptors has positive effects in computing correspondences between surfaces. As mentioned in Sec. 4.1, the functional map that relates the LBO eigenbases of two volumetric shapes also relates their restriction to the surfaces. Consequently, the volumetric map can be used to infer a correspondence between surfaces.
@@
Given two surface meshes $\partial\mathcal{M},\partial\mathcal{N}$, we compute the tetrahedral meshes $\mathcal{M},\mathcal{N}$ representing their interior volume. We then compute the functional map $\mathbf{C}$ relating their eigenbases via $\Phi_{\mathcal{M}}\mathbf{C}=T_{\pi}(\Phi_{\mathcal{N}})$, and we can use that same functional map for defining the surface relation $(\Phi_{\mathcal{M}})_{\partial}\mathbf{C}=(T_{\pi}(\Phi_{\mathcal{N}}))_{\partial}$. The surface correspondence is eventually extracted via nearest neighbor search, $\mathrm{NN}(\Phi_{\partial\mathcal{M}}\mathbf{C},\Phi_{\partial\mathcal{N}})$.
5 Results and Applications
We implemented our tool in Matlab, extending the surface only implementation of the functional map from [86] and ZoomOut [79]. Our code is available at https://github.com/filthynobleman/vol-fmaps.
5.1 Volumetric Functional Maps